OpenStudy (anonymous):

Help? I cant get this one... Choose the correct simplification of the expression (3x − 6)(2x^2 − 4x − 5). THANKS!

hero (hero):

Okay, hang on

hero (hero):

\[x = -\frac{\sqrt{14}-2}{2}, x = 2, x = \frac{\sqrt{14}-2}{2} \]

hero (hero):

Do you have a list of answer choices? Best to post them because I don't know how what form they want it simplified

OpenStudy (anonymous):

6x^3 -12x^2 -15x - 12x^2 + 24x +30 6x^3 -24x^2 + 9x + 30

hero (hero):

If I expand I get \[6x^3 - 24x^2 + 9x + 30\]

OpenStudy (anonymous):

Answer Choices 6x^3 − 24x^2 + 9x − 30 6x^3 + 9x + 30 6x^3 − 24x^2 + 9x + 30 6x^3 − 24x^2 + 39x + 30

hero (hero):

6x^3 − 24x^2 + 9x + 30 is the answer